OpenStudy (anonymous):

Solve using law of sines A = 76°, a = 15, b = 5

OpenStudy (anonymous):

want to try it 1st?\[\frac{ \sin A }{ a }=\frac{ \sin B }{ b }\]

OpenStudy (anonymous):

hey i'm tyring to do this one but since they gave me only one degree i'm having problems

OpenStudy (anonymous):

so far i got this \[15/\sin76 = 5/4.85 = c/sinC\]

OpenStudy (anonymous):

\[\frac{ sin 76 }{ 15 }=\frac{ sin B }{ 5 }\] solve for B

OpenStudy (amistre64):

gonna need an inverse tham

OpenStudy (anonymous):

angle C is 180-A-B

OpenStudy (anonymous):

so 104?

OpenStudy (anonymous):

but how can i find C when i only have one degree? 76?

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

B = 18.87

OpenStudy (anonymous):

C = 180 - 18.87 - 76

OpenStudy (anonymous):

how did you get 18.87?

OpenStudy (anonymous):

sin(76)/15=sin(B)/5 sin(B)=5*sin(76)/15 B=sin^-1(5*sin(76)/15)

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

so the degree of C will be 85?

OpenStudy (anonymous):

correct 85.13

OpenStudy (anonymous):

now find c

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so c is 4.98?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it should be 15.4

OpenStudy (anonymous):

but how ?

OpenStudy (anonymous):

hold on i think ididnt find what is sinB

OpenStudy (anonymous):

(15*sin(85))/sin(76)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

A = 76° B = 18.87° C = 85.13° a = 15 m b = 5 m c = 15.4 m

OpenStudy (anonymous):

ok i'm going to work it out

OpenStudy (anonymous):

got it! thank you!! :)